You have 1200 feet of fencing to enclose a rectangular garden. You want the width to be 20 feet less than the length. Find the dimensions of the garden.
Answer by Charles3475(23)

L = Length
W = Width = L-20 (given)
P = Perimeter = 2*L + 2*W (definition of perimeter)
P = 2*L + 2*(L-20) (substitution)
P = 1200 feet (given)
2*L + 2*(L-20) = 1200 (transitive property of equality)
2*L + 2*L - 2*20 = 1200 (distribution)
4*L - 2*20 = 1200 (addition of like terms)
4*L - 40 = 1200 (multiplication)
4*L = 1240 (add 40 to both sides)
L = 1240/4 (divide both sides by 4)
L = 310 (simplify)
W = L - 20 = 310 - 20 (substitution)
W = 290 (subtraction)
Happy Trails